OpenStudy (anonymous):

x and y vary inversely, and y = 3 when x = 6. Find y when x = 9. A. 2 B. 4.5 C. 12 D. 18

OpenStudy (jdoe0001):

"something" varies inversely to "something else" means \(\bf \textit{something}=\cfrac{\textit{some value}}{\textit{something else}}\implies y = \cfrac{\textit{some value}}{x}\) what's "n"? we dunno but we know that when x = 6, y =3, thus \(\bf y = \cfrac{n}{x}\implies 3 = \cfrac{n}{6}\implies 3\cdot 6=n\implies \color{blue}{y = \cfrac{3\cdot 6}{x}}\) what's "y" when x = 9? well \(\bf y = \cfrac{3\cdot 6}{9}\)

OpenStudy (jdoe0001):

\(\bf \textit{something}=\cfrac{\textit{some value}}{\textit{something else}}\implies y = \cfrac{\textit{some value}}{x}\implies y=\cfrac{n}{x}\) that is

OpenStudy (anonymous):

Wdf idk